Company XYZ is considering a project with the following projected cash flows:

CF0: -27,139

C01: 9,834

C02: 9,832

C03: 13,857

C04: 1,700

Calculate the Profitability Index (PI) of the project. Assume a WACC (discount rate) of 10.6%.

Enter your answer as a number with two decimal places of precision (i.e. 1.23).
Answer #1

Profitability index = Present value of cash inflows / initial investment

Present value of cash inflows = 9834 / (1 + 0.106)1 + 9,832 / (1 + 0.106)2 + 13,857 / (1 + 0.106)3 + 1,700 / (1 + 0.106)4

Present value of cash inflows = 8.891.500904 + 8,037.696732 + 10,242.44961 + 1,136.131009

Present value of cash inflows = $28,307.77826

Profitability index = 28,307.77826 / 27,139

Profitability index = 1.04